OpenStudy (anonymous):

when mulitply (sqrt8 -sqrt6)( sqrt8 + sqrt6) would this answer be right 1.994? simplify if possible????

jimthompson5910 (jim_thompson5910):

The answer is actually 2

OpenStudy (anonymous):

how did you get two

OpenStudy (anonymous):

You should use the algebraic identity: (a + b)(a - b) = a^2 - b^2.

jimthompson5910 (jim_thompson5910):

Use the difference of squares formula to find that \[\Large (\sqrt{8}-\sqrt{6})(\sqrt{8}+\sqrt{6})=(\sqrt{8})^2-(\sqrt{6})^2=8-6=2\] So \[\Large (\sqrt{8}-\sqrt{6})(\sqrt{8}+\sqrt{6})=2\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

i just realized you round it ,sorry am i right though? was my answer at least close?

jimthompson5910 (jim_thompson5910):

You were close but I'm not sure where you went wrong. Rounding won't always work

OpenStudy (anonymous):

No, you didn't round it, the answer is two.

OpenStudy (anonymous):

M-E-D-A-L?

OpenStudy (anonymous):

thanks..........

OpenStudy (anonymous):

thank you im looking at where i went wrong i think its when i times the two sqrts when i sqrt them i got the answer and that i think i dont know , but thanks again